OpenStudy (anonymous):

What method(s) would you choose to solve the equation: x2 + 2x - 6 = 0

OpenStudy (anonymous):

A. Square roots; there is no x-term. B. Quadratic formula, graphing; the equation cannot be factored easily since the numbers are large. C. Factoring; the equation is easily factored. D. Quadratic formula, completing the square or graphing; the coefficient of x2-term is 1, but the equation cannot be factored.

OpenStudy (anonymous):

i put b but it was wrong

OpenStudy (anonymous):

what does x2 mean? is it mean x^2

OpenStudy (anonymous):

i would say d ?

OpenStudy (anonymous):

yours must be \[x^2+ 2x-6 =0\], right?

OpenStudy (anonymous):

yes X^2

OpenStudy (anonymous):

ok. do you want to know how to get or just want the answer?

OpenStudy (anonymous):

i wanna know how to get the answer i know its b or d but not sure ?

OpenStudy (anonymous):

well its not b now

OpenStudy (anonymous):

ok, good. if so, i can help. tell me which numbers times each other to get 6

OpenStudy (anonymous):

you have so many options 6 *1 =6 3*2 =6 , right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i say its d

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

ok, if you think so, tell me the reason why you pick d

OpenStudy (anonymous):

if you have any idea, speak out. from then I can check or guide you if you stuck.

OpenStudy (anonymous):

ill simplify the work and give the answer and you tell me 3x * 2- 6 = 0 +6 +6 3x * 2 = 6 /2 /2 3x =3 /3 /3 x=1 does this help

OpenStudy (anonymous):

what do you think hoa

OpenStudy (anonymous):

I have no idea about your equation, from where? how you get 3x?

OpenStudy (anonymous):

i simplified x * 2 + 2x - = 0

OpenStudy (anonymous):

the formula for quadratic equation is x = \[\frac{ -b +(-)\sqrt{b^2 -4ac} }{ 2a }\]

OpenStudy (anonymous):

I asked the asker above, it's x^2 not 2x

OpenStudy (anonymous):

oh hehe fail for me i guess but which does my show

OpenStudy (anonymous):

ok, everyone has mistake. @hinarauf80 , your equation has a =1, b= 2 c = -6. replace all to the formula to get the answer. don't forget 1st is + the sqr, 2nd is - sqr. hope this help. if you cannot do it or stuck, let me know

OpenStudy (anonymous):

yea i think im lost know whats ^

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

oh well i tried see ya

OpenStudy (anonymous):

its ok just going over my test i got 4 wrong i am correcting it thank you anyway